OpenStudy (anonymous):

Use the line graph to solve. What was the difference in water levels between Week 2 and Week 4? A. 10 m B. 1 m C. 0.10 m D. 0.01 m

OpenStudy (anonymous):

line graph?

OpenStudy (anonymous):

is that all info?

OpenStudy (anonymous):

://static.k12.com/calms_media/media/713500_714000/713575/2/9dfdbaec1dd0e9a05eff595ddf0f512247bf7093/74592_question.jpg

OpenStudy (anonymous):

the total difference was .2m so dividing it by 2 i get the answer of 0.01m

OpenStudy (anonymous):

Thanks can you help me with 4 more

OpenStudy (anonymous):

can i see first

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Use the bar graph to solve. What fraction of the visitors are from Ohio or Michigan? A. B. C. D.

OpenStudy (anonymous):

the options please

OpenStudy (anonymous):

20+4+18+16) = total 20 + 16 = ohio + michigan 36/58= 18/29

OpenStudy (anonymous):

Lol Sorry I didnt notice that they didnt come out ok so A. 18/29

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Please post your questions on separate posts.

OpenStudy (anonymous):

B.19/29

OpenStudy (anonymous):

C. 5/9

OpenStudy (anonymous):

D. 4/5

OpenStudy (anonymous):

Whats the answer for this one though?

OpenStudy (anonymous):

18/29

OpenStudy (anonymous):

Ok thanks!

OpenStudy (anonymous):

Use the line plot to answer the question. How many people read two or more books during the summer? A. 6 B. 7 C. 13 D. 20

OpenStudy (anonymous):

Hold on i have to close this one and post the questions again